OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3.

OpenStudy (anonymous):

y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1 answers

OpenStudy (primeralph):